Seminar on mirror symmetry and representation theory

Tuesday, March 7, 5:00-6:30, MIT, room 2-132.

Quantization in positive characteristic and tilting of t-structures.

by Roman Bezrukavnikov (MIT).

In my first talk I have explained how tilting of t-structures and (affine) braid group actions arise in Bridgeland’s theory. After making some minor

corrections to what I have said then, I will explain how the method of quantization in positive characteristic (which comes from my joint work

with Mirkovic and Rumynin on geometric approach to modular representations) yields the same type of structures.

seminar contacts: Roman Bezrukavnikov (bezrukav@math.mit.edu); Rina Anno (rina@math.harvard.edu)